OpenStudy (anonymous):

The second term in a geometric sequence is 20. The fourth term in the same sequence is 45/4, or 11.25. What is the common ratio in this sequence?

OpenStudy (campbell_st):

ok... so you know the formula for a term in a geometric sequence \[a_{n} = a \times r^{n -1}\] a is the 1st term and r is the common ratio. so in your question you know 2nd term \[20 = ar^{2 -1}...or..... 20 = ar\] and the 4th term \[11.25 = ar^{4 -1}....or....11.25 = ar^3\] the 4th term can be written as ar times r^2 since r * r^2 = r^3 so you can substitute then 2nd term \[11.25 = ar \times r^2.... or........11.25 = 20 \times r^2\] now you just need to solve for r.

OpenStudy (anonymous):

would r be .75?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

thats right